# Asymmetric dynamical localization and precision measurement of BEC micromotion

### v1

@@ IV.0.2 Case II: optical lattice moving in lab frame
In this section, we discuss moving the lattice in the lab frame by inducing a constant frequency difference $\alpha$ in the range of $0-75$ kHz, which spans almost $10$ recoil momenta. After creating the BEC, $100\mu s$ time-of-flight is allowed, and subsequently periodic kicks are applied. As seen in Figs. 2(a-b), dynamical localization is observed for various lattice velocities.
Figure 2: Dynamically localization for two distinct frequency difference $\alpha$: (a) 15 kHz, (b) 30 kHz. (Top panel) experimental data shown as symbols for three different kick numbers, (bottom panel) simulation results shown as symbols at 70th kick.
For a direct comparison with case-I, optical lattice is moved with frequency difference $15$ kHz and $30$ kHz in the left direction, corresponding to $2\hbar k$ and $4\hbar k$ recoil momenta. The resulting steady-state momentum distribution after $50$, $60$ and $70$ kicks is displayed in Figs. 2 (a-b). The asymmetry induced by the lattice motion is visible for $15$ kHz and $30$ kHz; for $15$ kHz in Fig. 2(a) wave packet is moving in the same direction as a lattice (left direction), and for $30$ kHz in Fig. 2(b), the wave packet is moving in the opposite direction of the lattice (left direction).
The bottom panel in Fig. 2 shows the corresponding localization pattern obtained from QKR simulations. The simulation results confirm the emergence of asymmetric dynamical localization when the optical lattice is moved by inducing a frequency difference of $\alpha=15$ kHz and $\alpha=30$ kHz.
@@ V Measurement of early time dynamics
Figure 3: Oscillatory behaviour of $\langle p\rangle$ as a function of frequency difference (lattice velocity). In this experiment, $\langle p\rangle$ is measured after two kicks.
To understand asymmetry induced by relative motion between the atomic cloud and the optical lattice, early-time dynamics after two kicks is analyzed. The optical lattice is moved by creating a frequency difference between two laser beams, ranging from 0 to 75 kHz. Subsequently, two kicks are applied separated by time interval $T=24.3$ $\mu$s for different initial momenta $p_{0}$ and $\langle p\rangle$ is measured after a $10$ ms time-of-flight. Remarkably, pronounced oscillations in $\langle p\rangle$ are observed, consistent with Eq. 4. This is shown in Fig. 3, and suggests that $\langle p\rangle$ exhibits a linear relationship in the limit of $\alpha\to 0$. This linearity provides a promising avenue to measure the micromotion discussed in Section VI. Figure 3 also shows simulation results (blue line), which agrees with the experimental results in the limit $\alpha\to 0$. For the large $\alpha$, where the lattice velocity is high, experiment deviates from the simulation due to finite pulse time.
Another intriguing feature observed in Fig. 3 is the difference in the sign of $\langle p\rangle$ for $15$ kHz (corresponding to two-recoil momentum) and $30$ kHz (corresponding to four-recoil momentum). The opposite signs for $\langle p\rangle$ in the early-time dynamics is a consequence of the asymmetric distribution observed after long-time evolution, as demonstrated in Section IV. In general, short time asymptotic carry the signature of the long-term behavior at other frequencies as well. In particular, for frequencies that are integer multiples of 21 kHz, asymmetry is absent in the distribution. This corresponds to the condition $\alpha T=n/2$. Irrespective of the speed of lattice, if the frequency difference is an integer multiple of 20.57 kHz, no asymmetry is expected to manifest in the distribution, as discussed in Eq. 4.
@@ VI Measurement of micromotion of BEC
Figure 4: Measured $\langle p\rangle$ as a function of frequency difference $\alpha$ for two different time delays. Micromotion of BEC is measured by scanning over a range of frequency differences. See text for details.
Measuring the micromotion of the BEC poses a challenge due to its very small magnitude in the direction of the lattice [43]. This small magnitude of the velocity is difficult to measure accurately using conventional time-of-flight methods, which typically require long time-of-flight duration. One such measurement is performed in Ref. [21] through a $225$ ms time-of-flight. To address this challenge and accurately measure micromotion, early-time measurements are conducted by systematically tuning $\alpha$ of the lattice. A single-shot measurement can also measure micromotion if the lattice phase is stable. In these measurements, two different time delays, $T=24.3$ $\mu$s and $T=50$ $\mu$s, between two kicks, are employed.
The BEC micromotion is measured by tuning the lattice velocity in steps of frequency difference of $250$ Hz, as illustrated in Fig. 4. The phase evolution induced by micromotion is effectively balanced by the lattice motion. When the net average momentum $\langle p\rangle=0$, the BEC micromotion velocity can be deduced from the corresponding frequency difference. The calculated frequency difference at points of $\langle p\rangle=0$ are $(590\pm 45)$ Hz, corresponding to a velocity of $(460\pm 35)$ $\mu$m/s for $T=24.3$ $\mu$s, and $(649\pm 24)$ Hz, corresponding to a velocity of $(506\pm 19)$ $\mu$m/s for $T=50$ $\mu$s. The micromotion can also obtained by measuring asymmetry, keeping the frequency difference at $\alpha=0$, if the proportionality constant $c$ of Eq. 4 is known. However, former method is highlighted as it offers a direct and precise measurement of the micromotion.
@@ VII Conclusion
This work gives insights about the nature of dynamical localization when a wavefunction is launched with initial momenta $p_{0}\neq 0$ in the lab frame as well as when the optical lattice is moved in lab frame. In both scenarios, after a short diffusive timescale, the wavefunction is localized with an asymmetric distribution profile. This asymmetry emerges during the early time dynamics – driven by the breaking of parity symmetry due to the motion of the wavefunction or lattice. This feature is employed for precisely measuring the micromotion of the BEC. For the parameters of the experimental system we employed, velocity measurement yielded $(460\pm 35)$ $\mu$m/s in our system. This micromotion measurement is crucial for precision instruments such as atom interferometers and atomic gyroscopes, as micromotion can introduce systematic shifts and uncertainties in their measurements. The micromotion velocity is an order of magnitude smaller than the one recoil photon momentum, as well as mean velocity associated with BEC temperature. By utilizing the broken parity symmetry due to micromotion, measurement of such a small velocity is possible. Further, it might not be significantly affected by velocity distribution of the BEC, a common challenge in spectroscopy technique. This work contributes to our understanding of the precision measurement based on broken parity symmetry.
@@ VIII Acknowledgments
S.S.M. acknowledge a research fellowship from Council of Scientific and Industrial Research (CSIR), Government of India. All the authors thank the National Mission on Interdisciplinary Cyber Physical Systems for funding from the DST, Government of India through the I-HUB Quantum Technology Foundation, IISER-Pune.

### v3

@@ IV.0.2 Case II: optical lattice moving in lab frame
In this section, we discuss the dynamical localization by moving the lattice in the lab frame by inducing a constant frequency difference $\alpha$ in the range of $0-75$ kHz, which spans almost $5$ recoil velocity. After creating the BEC, $100\mu s$ time-of-flight is allowed, and subsequently periodic kicks are applied. As seen in Figs. 3(a-c), dynamical localization is observed for various lattice velocities.
Figure 3: Dynamically localization for three different lattice velocity: (a) $\hbar k/M_{Rb}$, (b) $2\hbar k/M_{Rb}$ and (c) $4\hbar k/M_{Rb}$. (left side) experimental data shown as symbols for three different kick numbers, (right side) simulation results shown as symbols at 70th kick.
For a direct comparison with case-I, the lattice has been moved in same direction, the observed distribution in this case has been flipped as discussed in Eq. 4 . The optical lattice is moved with velocity $\hbar k/M_{Rb}$ and $2\hbar k/M_{Rb}$ and $4\hbar k/M_{Rb}$ in the left direction. The resulting steady-state momentum distribution after $50$, $60$ and $70$ kicks is displayed in Figs. 3 (a-c). The asymmetry induced by the lattice motion is visible for $\hbar k/M_{Rb}$, $2\hbar k/M_{Rb}$ and $4\hbar k/M_{Rb}$; for $\hbar k/M_{Rb}$ in Fig. 3(a) the asymmetric wave packets are moving in the opposite direction as lattice (left direction), for $2\hbar k/M_{Rb}$ in Fig. 3(b), the asymmetric wave packets are moving in the same direction as lattice (left direction), and for $4\hbar k/M_{Rb}$ in Fig. 3(c), small asymmetric wave packet is moving again in the opposite direction as lattice.
The right panel in Fig. 3 shows the corresponding localization pattern obtained from QKR simulations. The simulation results confirm the emergence of asymmetric dynamical localization when the optical lattice is moved by velocities of $\hbar k/M_{Rb}$, $2\hbar k/M_{Rb}$ and $4\hbar k/M_{Rb}$.
@@ V Measurement of early time dynamics
To understand asymmetry induced by relative motion between the atomic cloud and the optical lattice, early-time dynamics after two kicks is analyzed. The optical lattice is moved by creating a frequency difference between two laser beams, ranging from 0 to $5\hbar k/M_{Rb}$. Subsequently, two kicks are applied separated by time interval $T=24.3$ $\mu$s for different initial velocity $v$ of lattice and $\langle p\rangle$ is measured after a $10$ ms time-of-flight. Remarkably, pronounced oscillations in $\langle p\rangle$ are observed, consistent with Eq. 4. This is shown in Fig. 4, and suggests that $\langle p\rangle$ exhibits a linear relationship in the limit of $v\to 0$. This linearity provides a promising avenue to measure the micromotion discussed in Section VI. Fig. 4 also shows simulation results (blue line), which agrees with the experimental results in the limit $v\to 0$. For the large $v$, where the lattice velocity is high, experiment deviates from the simulation due to finite pulse time.
Figure 4: Oscillatory behavior of $\langle p\rangle$ as a function of lattice velocity. In this experiment, $\langle p\rangle$ is measured after two kicks
Another intriguing feature observed in Fig. 4 is the difference in the sign of $\langle p\rangle$ for $\hbar k/M_{Rb}$ and $2\hbar k/M_{Rb}$. The opposite signs for $\langle p\rangle$ in the early-time dynamics is a consequence of the asymmetric distribution observed after long-time evolution, as demonstrated in Section IV. In general, short time asymptotic carry the signature of the long-term behavior at other velocities as well. In particular, for velocities that are integer multiples of $1.37\hbar k/M_{Rb}$ (observed $1.35\hbar k/M_{Rb}$ from Fig. 4), asymmetry is absent in the distribution. This corresponds to the condition $vT/\lambda=n/4$, when asymmetry is expected to vanish. Irrespective of the speed of lattice, if the velocity is an integer multiple of $1.37\hbar k/M_{Rb}$, no asymmetry is expected to manifest in the distribution, as a consequence of Eq. 4.
@@ VI Measurement of micromotion of BEC
Measuring the micromotion of the BEC poses a challenge due to its very small magnitude in the direction of the lattice [46]. This small magnitude of the velocity is difficult to measure accurately using conventional time-of-flight methods, which typically require long time-of-flight duration. One such measurement is performed in Ref. [23] through a $225$ ms time-of-flight. To address this challenge and accurately measure micromotion, early-time measurements are conducted by systematically tuning $\alpha$ (frequency difference) of the lattice. A single-shot measurement can also measure micromotion if the lattice phase is stable. In these measurements, two different time delays, $T=24.3\mu$s and $T=50\mu$s, between two kicks, are employed.
Figure 5: Measured $\langle p\rangle$ as a function of lattice velocity by changing the frequency difference $\alpha$ for two different time delays. Micromotion of BEC is measured by scanning over a range of lattice velocity. See text for details.
The BEC micromotion is measured by tuning the lattice velocity in steps of $0.017\hbar k/M_{Rb}$ corresponding frequency difference of $250$ Hz, as illustrated in Fig. 5. The phase evolution induced by micromotion is effectively balanced by the lattice motion. When the net average momentum $\langle p\rangle=0$, the BEC micromotion velocity can be deduced from the corresponding lattice velocity by linear fit. The calculated lattice velocity at points of $\langle p\rangle=0$ are $(0.039\pm 0.003)\hbar k/M_{Rb}$, corresponding to BEC micromotion velocity of $(230\pm 17)$ $\mu$m/s for $T=24.3$ $\mu$s, and $(0.043\pm 0.002)\hbar k/M_{Rb}$, corresponding to a velocity of $(254\pm 10)$ $\mu$m/s for $T=50$ $\mu$s. The micromotion can also be obtained by measuring asymmetry, keeping the lattice velocity at $v=0$, if the proportionality constant $c$ of Eq. 4 is known. However, the former method is favoured as it offers a direct and precise measurement of the micromotion.
@@ VII Conclusion
This work gives insights about the nature of dynamical localization under two scenarios; (i) when a wave packet is launched with initial momenta $p_{0}\neq 0$ in the lab frame, and (ii) when the optical lattice is moved in lab frame. In both scenarios, after a short diffusive timescale, the wavepacket is localized in momentum space with an asymmetric distribution profile. This asymmetry emerges during the early time dynamics – driven by the breaking of parity symmetry due to the motion of the wavepacket or the lattice. This feature is employed for precisely measuring the micromotion of the BEC. For the parameters of the experimental system we employed, velocity measurements yielded $(230\pm 17)$ $\mu$m/s for $T=24.3$ $\mu$s and $(254\pm 10)$ $\mu$m/s for $T=50$ $\mu$s in our system. This micromotion measurement is crucial for precision instruments such as atom interferometers and atomic gyroscopes, to correct for systematic shifts and uncertainties. The micromotion velocity is an order of magnitude smaller than one recoil photon momentum, as well as comparable to mean velocity associated with BEC temperature. The broken parity symmetry induced by the micromotion is utilized to measure such small velocity. Further, it might not be significantly affected by velocity distribution of the BEC, a common challenge in spectroscopy technique. This work contributes to our understanding of the precision measurement techniques using BEC-based quantum kicked rotor models.
@@ VIII Acknowledgments
S.S.M. acknowledges research fellowship from Council of Scientific and Industrial Research (CSIR), Government of India. All the authors thank the National Mission on Interdisciplinary Cyber Physical Systems for funding from the DST, Government of India through the I-HUB Quantum Technology Foundation, IISER Pune.